# Against The Achilles’ Heel: A Survey on Red Teaming for Generative Models

### v1

@@ 4.2.2 Search Goal
Searchers with direct goals focus on finding prompts that induce dangerous content. However, evaluating the dangerous content is a nuanced topic. Common evaluators are either based on fine-tuned language models (Casper et al., 2023; Wichers et al., 2024) or meticulously prompted large language models like GPT4 (Chao et al., 2023; Mehrotra et al., 2023; Jin et al., 2024). Although direct goals are conceptually straightforward, there are several problems to consider. Classifiers contain biases rooted in their training data distribution, which limits the searcher exploration space and may be exploited in certain cases, resulting in inferior search results. There is a trade-off between iteration speed and accuracy, as larger classifiers are generally more accurate but also slow to perform inference.
@@
The selection of proxy goals limits the search space and accelerates search progress. A large number of works focus on instruction indirection (as mentioned in Sec. 4.1.2), that is, increasing the linguistic complexity of the prompt to hide the true intention (Zhang et al., 2023c; Yao et al., 2023a; Jiang et al., 2023d). The problem with this proxy goal is that the correlation between complexity and attack success varies considerably with language models.
Another branch exploits completion compliance (as discussed in Sec. 4.1.1) combined with optimization algorithms, especially the affirmative suffixes, with notable examples of AutoDAN-I (Zhu et al., 2023). The primary objective of AutoDAN-I’s optimization is to minimize the likelihood of generating target harmful statements, achieving this through adjustments in adversarial suffixes. The AutoDAN-I and PAL (Sitawarin et al., 2024) use the affirmative suffix strategy (refer to Sec. 4.1.1) as the main optimization goal: they try to find a suffix such that, once appended to the prompt, the language model starts their response with affirmative suffixes like “Sure, here is how”. This shrinks the search complexity considerably, and the suffix is evasive to the filter. For better interpretability, AutoDAN-I adds a loss term to ensure low perplexity of the suffix, and PAL uses a proxy model to filter for interpretable suffixes. TrojLLM (Xue et al., 2023) combines the complexity with reinforcement learning, with more details discussed in Sec. 4.2.3.
@@ 4.2.3 Search Operation
We classify search operations based on their optimization formulation: continuous and discrete. Continuous search operations rely on gradients to guide the optimization, such as Proximal Policy Optimization (PPO) or Greedy Coordinate Gradient (GCG). Discrete search directly operates with prompts or suffixes on themselves, such as genetic algorithms and language model rewriting. Continuous search operations need to tackle the problem of approximating the gradient as language is naturally discrete.
@@
  • Reinforcement Learning The RL agent interacts with a language model by submitting prompts or suffixes and receives feedback based on the effectiveness of these submissions. For example, Xue et al. (2023) formulate the problem as identifying an optimal combination of trigger and prompt, adopting a two-stage approach of first finding a prompt seed that leads to high accuracy, and then using an agent to search a trigger token that maximizes a reward function. This dual-focus approach underscores the RL agent’s ability to adapt and learn from feedback, optimizing strategies that navigate the delicate balance between operational integrity and the effectiveness of the attack. Casper et al. (2023) uses Proximal Policy Optimization(PPO) to fine-tune GPT-2-large to produce harmful prompts deemed by the classifier. Their reward function contains the classifier’s logit confidence and the diversity that penalizes the similarity of prompts measured by the cosine distance of the target LM’s embedding of prompts.
  • Other Approaches include the COLD-Attack (Guo et al., 2024) where they apply the COLD, a controllable text generation method, to achieve better control over the attack direction. In this method, the attack constraints, including the affirmativeness, fluency, and lexical constraints, are expressed with a compositional energy function that modifies the vanilla logits of language models with Langevin Dynamics before decoding.
@@
  • Concatenation is putting different templates together to increase Instruction Indirection (according to Sec. 4.1.2). This is adopted by a series of studies for its simplicity: PromptPacker (Jiang et al., 2023d), FuzzLLM (Yao et al., 2023a), Puzzler (Chang et al., 2024) among others. Jin et al. (2024) proposed a random walk-based method on knowledge graphs built from existing jailbreaks to explore jailbreak fragments.
@@
  • Genetic Algorithm The Genetic Algorithm adds more diversity by employing mechanisms like crossover, which merges elements from two different prompts, and mutation, which introduces random variations into a prompt. These operations are designed to select and propagate combinations with the greatest potential for achieving the objective across successive generations of prompts. This branch of composition is less explored, with Lapid et al. (2023) and AutoDAN-GA (Liu et al., 2023d) being the exception.
  • Language Model This method relies on the language model’s ability to understand the context and intent of the original prompt while generating new prompts that might lead to the desired jailbreak behavior. The challenge here lies in guiding the model to produce variations significantly different from the original while aiming for the same jailbreak goal. To ensure the effectiveness of using language models in this scenario, previous work has employed both general instructions such as “acting as a helpful red-teaming assistant” (Takemoto, 2024a; Chao et al., 2023; Mehrotra et al., 2023; Takemoto, 2024a; Deng et al., 2023a), and specific instructions on the use of certain strategies (Jiang et al., 2023d; Liu et al., 2023a).

### v2

@@ 4.2.2 Search Goal
Searchers with direct goals focus on finding prompts that induce dangerous content. However, evaluating the dangerous content is a nuanced topic. Common evaluators are either based on fine-tuned language models Casper et al. (2023); Wichers et al. (2024) or meticulously prompted LLMs like GPT4 Chao et al. (2023); Mehrotra et al. (2023); Jin et al. (2024). Although direct goals are conceptually straightforward, there are several problems to consider. Classifiers contain biases rooted in their training data distribution, which limits the searcher exploration space and may be exploited in certain cases, resulting in inferior search results. There is a trade-off between iteration speed and accuracy, as larger classifiers are generally more accurate but also slow to perform inference.
@@
The selection of proxy goals limits the search space and accelerates search progress. A substantial body of work focuses on instruction indirection (as mentioned in Section 4.1.2), that is, increasing the linguistic complexity of the prompt to hide the true intention Zhang et al. (2023c); Yao et al. (2023a); Jiang et al. (2023d). The problem with this proxy goal is that the correlation between complexity and attack success varies considerably with language models.
Another branch exploits completion compliance (as discussed in Section 4.1.1) combined with optimization algorithms, especially the affirmative suffixes, with notable examples of AutoDAN-I Zhu et al. (2023). The primary objective of AutoDAN-I’s optimization is to minimize the likelihood of generating target harmful statements, achieving this through adjustments in adversarial suffixes. AutoDAN-I and PAL Sitawarin et al. (2024) use the affirmative suffix strategy (refer to Section 4.1.1) as the main optimization goal: they try to find a suffix such that, once appended to the prompt, the language model starts the response with affirmative suffixes like “Sure, here is how”. This shrinks the search complexity considerably, and the suffix is evasive to the filter. For better interpretability, AutoDAN-I adds a loss term to ensure low perplexity of the suffix, and PAL uses a proxy model to filter for interpretable suffixes. TrojLLM Xue et al. (2023) combines the complexity with reinforcement learning, with more details discussed in Section 4.2.3.
@@ 4.2.3 Search Operation
We classify search operations based on their optimization formulation: continuous and discrete. Continuous search operations Proximal Policy Optimization (PPO) or Greedy Coordinate Gradient (GCG). Discrete search directly operates with prompts or suffixes on themselves, such as genetic algorithms and language model rewriting. Continuous search operations need to tackle the problem of approximating the gradient as language is naturally discrete.
@@
  • Reinforcement Learning (RL) The RL agent interacts with a language model by submitting prompts or suffixes and receives feedback based on the effectiveness of these submissions. For example, Xue et al. (2023) formulate the problem as identifying an optimal combination of trigger and prompt, adopting a two-stage approach of first finding a prompt seed that leads to high accuracy, and then using an agent to search a trigger token that maximizes a reward function. This dual-focus approach underscores the RL agent’s ability to adapt and learn from feedback, optimizing strategies that navigate the delicate balance between operational integrity and the effectiveness of the attack. Casper et al. (2023) uses PPO to fine-tune GPT-2-large to produce harmful prompts deemed by the classifier. Their reward function contains the classifier’s logit confidence and the diversity that penalizes the similarity of prompts measured by the cosine distance of the target LM’s embedding of prompts.
  • Other Approaches include the COLD-Attack Guo et al. (2024) where they apply the COLD, a controllable text generation method, to achieve better control over the attack direction. In this method, the attack constraints, including affirmativeness, fluency, and lexical constraints, are expressed with a compositional energy function that modifies the vanilla logits of language models with Langevin Dynamics before decoding.
@@
  • Concatenation is putting different templates together to increase Instruction Indirection (see Section 4.1.2). This is adopted by a series of studies for its simplicity: PromptPacker Jiang et al. (2023d), FuzzLLM Yao et al. (2023a), Puzzler Chang et al. (2024) among others. Jin et al. (2024) proposed a random walk-based method on knowledge graphs built from existing jailbreaks to explore jailbreak fragments.
@@
  • Genetic Algorithms Genetic Algorithms add more diversity by employing mechanisms like crossover, which merges elements from two different prompts, and mutation, which introduces random variations into a prompt. These operations are designed to select and propagate combinations with the greatest potential for achieving the objective across successive generations of prompts. This branch of composition is less explored, with Lapid et al. (2023) and AutoDAN-GA Liu et al. (2023d) being rare examples of this approach.
  • Language Model Generation This approach uses languages models to generate new prompts based on past jailbreak attempts to lead to desired jailbreak behavior. In contrast to continuous approaches where language models perform search by updating continuously with gradients, Language Model Generations explore the state space by generating discrete tokens, though it may be argued that the implicit gradient updates underlie the in-context learning process von Oswald et al. (2024). To ensure the effectiveness of using language models in this scenario, previous work has employed both general instructions such as “acting as a helpful red-teaming assistant” Takemoto (2024a); Chao et al. (2023); Mehrotra et al. (2023); Takemoto (2024a); Deng et al. (2023a), and specific instructions on the use of certain strategies Jiang et al. (2023d); Liu et al. (2023a).
4.3 Analysis of Attack Methods
The landscape of LLM attack methods reveals complex interactions and trade-offs between different approaches.
4.3.1 Fundamental Trade-offs and Relationships
The field of LLM attacks is characterized by two primary approaches: manual strategy development and automated searching, each with distinct characteristics and implications. Manual strategies typically achieve higher attack diversity and better exploitation of model capabilities, demonstrating remarkable creativity in circumventing model safeguards. However, they lack systematic coverage and scalability, making it difficult to guarantee comprehensive testing of model vulnerabilities. Automated searching methods, conversely, offer systematic exploration and scalability but often produce less diverse attacks, potentially missing creative attack vectors that human researchers might discover. This fundamental tension between creativity and systematicity remains a central challenge in the field.
The relationship between different attack categories reveals important patterns in how vulnerabilities can be exploited. Completion compliance attacks often serve as a foundation upon which more sophisticated approaches are built. For instance, when attackers combine the affirmative suffix strategy with instruction indirection techniques, they create more robust attacks that are harder to defend against. This layering effect demonstrates how basic completion patterns can be enhanced through complex instructional frameworks, resulting in more effective attacks.
4.3.2 Comparative Analysis of Attack Methods
A detailed examination of different attack methods reveals varying effectiveness and applicability across different contexts. Completion compliance methods (Section 4.1.1) have shown high initial success rates but face increasing challenges as detection mechanisms improve. Their primary advantage lies in their simplicity and low implementation cost, making them attractive for initial testing. However, their effectiveness has declined as models incorporate better defenses against such basic attacks.
Instruction indirection methods (Section 4.1.2) represent a more sophisticated approach, demonstrating variable success rates but greater sustainability over time. Their superior adaptability across different contexts comes at the cost of more complex implementation requirements. The semantic complexity of these attacks presents significant challenges for detection mechanisms, as they often mirror legitimate usage patterns.
Generalization glide methods (Section 4.1.3) have emerged as particularly powerful attacks, achieving high success rates by exploiting fundamental model properties. Their effectiveness across languages and modalities makes them especially challenging to defend against, as they leverage the model’s inherent generalization capabilities rather than obvious vulnerabilities.
Model manipulation methods (Section 4.1.4), while requiring the highest level of technical expertise and resources, achieve remarkable success rates when applicable. Their effectiveness stems from direct intervention in the internals of the model, though their applicability is limited by access requirements and implementation complexity.
5 Safeguarding of Large Language Models
Given the diversity of attack types, defense measures have been widely researched and adopted in language model training and deployment. Defenses can be implemented in two ways: at training time and at inference time (Figure 20). We use training time to refer to the supervised fine-tuning and the reinforcement learning from human feedback (RLHF) step after pretraining. In this stage, model parameters are updated to improve a model’s ability to discern attacks and reject them, whereas at inference time, model weights are frozen and its behavior can only be controlled by prompting. Guardrails that filter out unsafe content are also common in inference-time defenses.
Figure 20: Defense strategies implemented at different stages in the life cycle of a language model. Training-time defense involves steps such as RLHF and fine-tuning to implement safety alignment, while inference-time defense employs system prompting, filtering out unsafe content, and guardrailing systems to direct model behavior.
These two types of defense have different merits and drawbacks. In comparison to training-time defense, inference-time defense is more flexible and modular, as prompts can be edited and components can be introduced to the pipeline independently. However, all these measures add latency to the system and may not be able to identify nuanced attacks or those that require more reasoning. Training-time defense provides models with the ability to recognize intricate attacks. However, these measures are shown to impose an alignment tax Lin et al. (2024) that impairs the normal function of language models, as the process shifts the distribution and causes catastrophic forgetting.
5.1 Training-time Defense
In this section, we discuss training-time defense, covering different approaches through the pipeline of pre-training and post-training.
5.1.1 Pre-training
Defense efforts begin as early as the pre-training stage. Some approaches, such as those proposed by Team et al. (2024); Dubey et al. (2024); AI et al. (2024), involve filtering harmful content from the pre-training corpus that violates strict safety policies. Additionally, some methods adopt conditional pre-training Anil et al. (2023), where special control tokens for toxicity labels assigned by classifiers, are added to the training data. This enables the model to learn more structured text representations during pre-training. Implementing safety measures at the pre-training stage can effectively reduce the challenges of safety alignment in the post-training phase Team et al. (2024).
5.1.2 Post-training
Supervised Fine-tuning (SFT)
The most basic way of supervised fine-tuning for defense is to fine-tune language models with samples that include unsafe prompts followed by rejections. Bianchi et al. (2023) discovered that fine-tuning the Llama model with just an additional 3% of such samples can notably enhance its security. Other work has looked into more sophisticated pipelines. Central to fine-tuning is the collection of rejection data, which needs to be diverse to enable generalization across different domains. Ge et al. (2023) collected fine-tuning data through an adversarial approach, where one adversarial model is trained to generate unsafe queries, while a second model is safety-tuned to refuse such queries. The fine-tuning data is selected through a helpfulness reward model and a safety reward model, where unsafe data rated by the safety reward model is used for adversarial training, while the safe and helpful data is used for safety fine-tuning. Wang et al. (2023h) first fine-tuned the language model to classify questions as harmful or harmless. Then, the same model is further fine-tuned to determine whether its responses are harmful. During inference, the language model will add a safety label to its responses, which can be used to filter out unsafe responses. On top of that, Madry (2017) proposed an adversarial training framework to train models that are resistant to adversarial attacks. Their approach involved formulating a min-max problem to enable security against adversarial perturbations, and demonstrated that with sufficient model capacity and training against strong adversaries like projected gradient descent (PGD), networks could be made significantly more robust against a wide range of attacks.
Considering the training cost for fine-tuning, there are also training-efficient defense methods that use soft-prompt tuning He & Vechev (2023) and LoRA Varshney et al. (2024).
Preference Alignment by RLHF
In addition to supervised fine-tuning, RLHF is widely used by both industry and academia to instill human preferences and safety measures into LLMs. In RLHF, a reward model is trained to learn human preferences and then used to tune the target language model. The dataset for tuning reward models is the rank of responses to the same questions rated by the labeler Ouyang et al. (2022). Compared to fine-tuning, RLHF has been shown to improve out-of-distribution (OOD) generalization, which is for crucial safeguards Kirk et al. (2023b).
A key problem in performing RLHF is the construction of the preference dataset. Ji et al. (2023) built a safety dataset for the question-answering task. In this dataset, pairs of answers with expert annotations of helpfulness and harmlessness are created for each question. Shi et al. (2023b) proposed a more scalable automatic approach for constructing preference datasets. Specifically, they created a reverse instruction model for generating instructions given specific texts. For example, given a sonnet praising the value of human love, the model would generate an a instruction such as “Write a sonnet around human love”. In their case, they used the model to generate questions from harmful content, forming question–answer pairs for preference datasets.
Preference Alignment by RLAIF
While previous work has focused on the annotation of safety data, existing datasets often fall short of addressing the full range of user-specific safety requirements, such as differences in values among diverse groups or the particular terms of service for a given organization. Developing new annotated datasets demands substantial human labor, though the creation of annotation rules themselves requires only minimal labeling effort. Since the work of  Bai et al. (2022b), numerous approaches have explored using models to assist in annotating alignment data based on a predefined set of safety rules, referred to as a “Constitution”. Similar to RLHF, this approach involves training a preference model to align model behavior to desired standards. However, in contrast to RLHF, it substitutes human feedback on harmlessness with “AI feedback”, and is thus termed RLAIF. RLAIF substantially reduces the human effort required for annotation and has been widely adopted in methods such as Claude Anthropic (2024), Gemini Team et al. (2024), and Qwen2 Yang et al. (2024a).
Preference Alignment Algorithms
Given preference data, how should model parameters be updated? In InstructGPT Ouyang et al. (2022), OpenAI used the proximal preference optimization (PPO) algorithm Schulman et al. (2017). This approach requires training a preference model based on preference data. The training process is dynamic: in each iteration, the current model generates outputs, which are then scored by the preference model, and these scores are used to update the model through gradient-based optimization. Since this method requires additional training of a preference model along with dynamic data generation and scoring, it offers highly tailored optimization for each model. However, it also adds complexity to the training process, making it challenging to implement. There are variants of PPO, for instance, RRHF Yuan et al. (2023a) which applies a ranking loss, and GRPO Shao et al. (2024) which optimizes the memory usage of PPO.
In comparison, the direct preference optimization (DPO) algorithm Rafailov et al. (2023) formulates the process as fine-tuning and updates the model’s parameters directly using preference data. DPO greatly simplifies the optimization process, resulting in more stable training than PPO. Consequently, DPO has been widely adopted by models including Zephyr Tunstall et al. (2024), Qwen2 Yang et al. (2024a), Llama3 Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Yi AI et al. (2024). Nonetheless, the feedback introduced in DPO is independent of a specific model, and its limited generalizability has been noted in several studies Xu et al. (2024a).
The preference optimization process needs to balance multiple objectives to meet human preferences. One pair of contrasting objectives is helpfulness and harmlessness. These objectives are often conflicting because it is desirable for models to comply with normal questions (reflecting their helpfulness), but refuse to answer harmful questions (providing harmlessness), which creates a delicate balance. Such multi-objective optimization is often unstable and prone to mode collapse, making the model lean one way or the other. To tackle this problem, Dai et al. (2023) separate these objectives during optimization by framing harmlessness as a cost objective and optimizing helpfulness with Lagrangian optimization, achieving a model with a better trade-off between helpfulness and harmlessness.
Beyond these optimization algorithms, other general preference alignment algorithms can support safety alignment Jiang et al. (2024b). As this paper is focused specifically on red teaming, these alternative methods will not be discussed further here.
Representation Engineering
Apart from updating models end-to-end with gradient-based approaches, a number of studies have tried to approach the problem of training-time defense in a more mechanistically interpretable way. To further improve controllability and reliability, they have explored directly analyzing and modifying the internal representations of models. Early work focused extensively on examining internal representations Zou et al. (2023a); Caron et al. (2021). For example, knowledge editing Mitchell et al. (2022); Meng et al. (2022); Xu et al. (2023f) tries to directly edit the MLP layer of transformer language models to erase or introduce new knowledge, and other steering methods take a similar approach Upchurch et al. (2017); Ilharco et al. (2023); Turner et al. (2023). Recent research has further investigated the feasibility of employing representation engineering as a defense mechanism Zou et al. (2023a); Li et al. (2024e; 2023c); Zou et al. (2024a); Yao et al. (2023d). For instance, Zou et al. (2024a) redirect the representations of potentially harmful responses to their orthogonal representations or EOS tokens, causing the model to generate null or interrupted responses in place of harmful ones.
Additionally, machine unlearning, a technique traditionally used to address privacy and copyright issues Nguyen et al. (2022); Yao et al. (2023d); Qu et al. (2023); Lu et al. (2024b), can also be employed to mitigate security concerns. For example, Yao et al. (2023d) enhanced model output safety by performing gradient ascent and random sampling on harmful data.
5.2 Inference-time Defense
Inference-time defense generally takes the form of different prompting techniques and guardrail systems. We also focus on emerging topics such as language model ensembles and adversarial suffix filters that target those generated by the AutoDAN searcher Zou et al. (2023b) (Table 2).
5.2.1 Prompting
Language models exhibit remarkable capabilities in in-context learning Brown et al. (2020) and following instructions Ouyang et al. (2022), paving the way for the use of prompting as a training-free approach to enhance the safety of language models. We classify such prompting methods into the following categories.
Prompt Rewriting
Rewriting the input prompt is a straightforward method to enhance response safety. Xie et al. (2023) add cues about responsibility and harmlessness directly to the system prompt, and in doing so improved the rejection rate of prompts designed to jailbreak or manipulate the model. RPO Zhou et al. (2024a) uses a similar approach to GCG Zou et al. (2023b): rather than searching for an adversarial suffix, it searches for a suffix that could reduce harm, after which models are less likely to respond with harmful content.
Figure 21 shows a defense method based on self-reflection. The prompt first asks the LLM to perform the task while following safety guidelines, then reflects and double-checks the outputs for safety and ethics.
Figure 21: A defense method based on self-reflection from  Wu et al. (2023a).
Few-shot Prompting
Furthermore, few-shot prompting has been extended to counteract sophisticated adversarial attacks. The strategy of In-Context Defense Wei et al. (2023b) integrates specific examples of attack refusals into the prompt, thereby improving the recognition and resilience of the model to various attacks. Figure 22 demonstrates the process. Given the constraints on context length and the diversity of user queries, the combination of retrieval and few-shot prompting Rubin et al. (2022) is often used to ensure more effective and tailored responses to a wide range of user queries Meade et al. (2023). This method retrieves safety-related examples that closely match the user’s query to be added to the context, providing a more comprehensive defense across different potential areas of harm.
Figure 22: An illustration of few-shot attack and defense methods (figure from Wei et al. (2023b)).
Chain-of-Thought
Adopting a Chain-of-Thought (COT) Wei et al. (2022) approach enhances reasoning of language models, which has been extended to discern complex attack patterns in several studies. Zhang et al. (2024c) use a multi-stage procedure that examines the intention behind a query before generating a response with established policies. Explicitly discerning the intent is particularly advantageous as language models often lack the ability to retrieve and reason on knowledge Allen-Zhu & Li (2023); Berglund et al. (2023) implicitly.
Compared to training-time defenses, prompting is a cost-effective method of implementing safety measures because it avoids modifying the LLM weights. However, additional system prompts might increase response latency, particularly when numerous examples are used or multi-stage prompting is involved. Additionally, including unsafe content within examples can inadvertently guide language models during response and pose a risk of exploitation. In practice, prompting is often used with other techniques like keyword filtering.
5.2.2 Guardrail Systems
To systematically control language model responses, guardrail systems have been developed to provide a unified interface to filter unsafe content Dong et al. (2024a), often through a domain-specific language. The NeMo Guardrails system Rebedea et al. (2023) provides a method to employ LLMs and vector databases to check unsafe content and hallucination in designated points of a dialogue flow, using the Colang language. Similarly,  Sharma et al. (2024) devised SPML, a domain-specific language that empowers prompt developers to efficiently create and maintain secure system prompts. SPML employs an intermediate representation for each entry, facilitating the comparison of incoming user inputs to ensure their safety.  Rai et al. (2024) built a multi-tiered guardrail system that includes a system prompt filter, toxic classifier, and ethical prompt generator to filter unsafe content,
To better identify unsafe content, fine-tuned models that check prompts for harmfulness have also been developed. Similarly, Pisano et al. (2023) added a prompt detecting stage and a response correcting stage where another model critiques incoming queries and the generated responses. The critique is appended to the query as extra signal to the primary model. Inan et al. (2023) fine-tuned a Llama2 7b model to classify prompts based on their proposed taxonomy and risk guidelines, and found the model to exhibit a high degree of transferability to other guidelines.
5.2.3 Language Model Ensembling
Language Model Ensembling refers to defense methods based on synthesizing and summarizing predictions from multiple models to derive answers. Chen et al. (2023a) introduced Moving Target Defense (MTD), which combines responses from eight commercial large-scale models to form the final answer. The authors devised their evaluation model to select “harmless” and “helpful” responses. While the ensembling method is conceptually simple, it inevitably leads to long response times and high computational costs. Hence, its practical utility is limited.
Chern et al. (2024) further extended the method to have multiple models debate amongst themselves. In each multi-agent debate session, given a sensitive or dangerous topic, each LLM agent is first prompted to give a zero sample initial response. Then, for the number of rounds specified by the user, the agents perform a “discussion”, in which each agent updates its response with the output of other LLM agents (or itself) as additional recommendations. The authors found the approach to enhance safety especially for smaller models when combined in debate with larger models, but again at a prohibitively high inference cost.
5.2.4 Against Adversarial Suffixes
It has been found that by appending certain suffixes to the harmful query, the model can be induced to start the response with affirmative phrases and respond with harmful content Zou et al. (2023b). This approach is difficult to detect with conventional keyword input filters. Several approaches have been proposed:
  • Perplexity Filter These defenses are based on the fact that the suffixes are often nonsensical and have high average perplexity Jain et al. (2023). However, Alon & Kamfonas (2023) found that relying solely on perplexity filtering could lead to a high false positive rate. To counteract the problem, they train a classifier using the Light Gradient Boosting Machine (LightGBM) algorithm. They then run predictions on all validation samples and tune the thresholds that map the results to positive (adversarial) or negative (non-adversarial) classes to maximize the $F_{1}$ score.
  • Perturbation Another feature of these suffixes is that they are very sensitive to small perturbations. Based on this observation, Kumar et al. (2023) developed the erase-and-check algorithm to detect and pre-process harmful prompts. Given a prompt $P$, this process removes the tokens one by one (up to $d$ tokens) and uses a security filter to check whether the deleted subsequence is safe. If the input prompt $P$ or any of its deleted subsequences are detected as harmful, the input prompt is marked as harmful. Instead, the prompt $P$ is marked as safe only if the filter detects all sequences that are checked as safe; see the example in Figure 23.
    Figure 23: Example of filter defense based on token erasure. The prompt can be correctly identified as harmful by LLMs after the harmless suffixes are striped in the token-erasure process.
    Robey et al. (2023) proposed the SmoothLLM algorithm based on randomized smoothing, a method to improve the adversarial robustness of models. It defends against attacks against samples by adding random noise to the input data. They enhanced the adversarial robustness of LLMs by employing insert, swap, and patch perturbations to their jailbreak attempts. This approach effectively reduces the success rate of GCG attacks to below 1%.
Figure 24: Defense strategies for corresponding attack methods. Most training-time and inference-time defense methods are universal against attacks. The suffix defense methods specifically mitigate the adversarial suffix methods introduced in Section 4.2.
5.3 Comparative Analysis of Defense Methods
Different defensive mechanisms exhibit varying applicability depending on their context. Training-time defenses are more adept at detecting subtle and multi-step attacks compared to inference-time defenses. However, they may alter the model’s distribution, potentially compromising the language model’s performance. Inference-time defenses, on the other hand, offer greater flexibility, allowing for adaptive adjustments in response to diverse attack strategies. Despite this adaptability, they are generally less effective against complex attacks.
Pre-training defenses (Section 5.1.1) effectively mitigate most conventional risk issues. In contrast, supervised fine-tuning, as described in post-training defense (Section 5.1.2), provides more nuanced protection for specific tasks and requires only a minimal amount of task-specific data, making it cost-efficient. Reinforcement Learning from Human Feedback (RLHF) significantly enhances the robustness of model security defenses, ensuring alignment with human expectations and requirements. Additionally, the introduction of Reinforcement Learning from AI Feedback (RLAIF) addresses the high costs associated with creating preference datasets.
Prompting (Section 5.2.1) is a cost-effective method that is moderately successful. However, it introduces latency and poses a security risk in which examples might be exploited. In contrast, guardrail systems (Section 5.2.2) enforce the filtering of unsafe content by incorporating explicit constraints and rules. While effective, these systems must be meticulously designed to avoid impinging on regular usage.
Language Model Ensembling (Section 5.2.3) offers high robustness and reliability, and provides more balanced and fair results. However, this method incurs higher costs due to the need to generate outputs from multiple models. Approaches to specifically address adversarial suffixes (Section 5.2.4) are effective against suffix-based attacks, but less effective against more diverse attacks.
Most of these defense methods provide broad-spectrum protection against the attacks discussed in this paper by adjusting prompts (for Prompting and Ensembling methods) or modifying training data (for Training-time Defense and Guardrail Systems) at the input/output level, as illustrated in Figure 24. However, certain methods are designed specifically to counter targeted attacks. The defense approach detailed in Section 5.2.4 is tailored specifically to protect against the suffix-based attack described in Section 4.2.